Q1: Using CLV to evaluate spending on customer acquisition and customer retention.

Brushes & Buckets, a paint wholesaler, sells paint to professional painters. On average, each painter spends $38,333 per year on paint. The average gross margin for Brushes & Buckets is 35%, and it currently has $4.2 million of sales in this segment. The current retention rate of each painter is 80%, and the revenue per painter is stable over time. (Assume a 10% discount rate.)

  1. The marketing director proposes spending $1,000,000 for a one-time direct marketing campaign to acquire new professional painters. How many painters would Brushes & Buckets need to acquire to make this investment profitable in the long term?   (10 points)

Hint: Use CLV to help calculate the breakeven point.

CLV = = m [ r / (1+i-r)]

m = gross margin

r = retention rate

i = discount rate

  1. The marketing director also wants to determine the maximum amount of money Brushes & Buckets should spend to increase the retention rate of the current professional painters from 80% to 90%.

Hint: Use CLV to help calculate the difference in Lifetime value when the retention rate is 80% vs. 90%.
Answer #1

Answer :

(a)

CLV (Customer life time value) = Gross margin * (Retention rate / [ 1 + Rate of discount - Retention rate ]

CLV (Customer life time value) = 38333*35% *(80% / (1+10%-80%)

CLV (Customer life time value) = 35,777.4667

Professional painter to be acquired for spent of 1,000,000 (b) = 27.9505 ((1,000,000/35,777.47)

(b).

CLV (Customer life time value) with retention rate 90% 38,333*35%*(90%/(1+10% - 90%)
CLV (Customer life time value) with retention on rate 90% (c) 60,374.48
Revenue (d) 4,200,000.00
Average spent by painter (e) 38,333.00
Number of painter on board (f= d/e) 109.57
Additional contribution due to increased retention rate per painter (g = c - a) 24,597.01
Total additional contribution due to increased retention rate (h = g*f) 2,695,000.00

Please Upvote Thank You !!!